A a straight horizontal bar of mass 10 mg and length 1m carries a current of 2 ampere what minimum magnetic field should be applied in the region so that the magnetic force on the wire may balance it's weight

Grade:

1 Answers

Susmita
5 years ago
Magnetic force on the wire F=IBL=2*B*1=2B
Weight=mg=10×10-3×10=0.1
F=0.1
Or,B=0.1/2=0.05T=50 mT
If it helps please approve.